Dan is preparing for a 10-mile race. He will run the first 5 miles at an average speed of 6 miles per hour, the next 3 miles at an average speed of 8 miles per hour, and the final 2 miles at an average speed of 10 miles per hour. How many hours, rounded to the nearest hundredth of an hour, will it take Dan to run the race if he follows his plan?
Answer by mananth(16949)

Dan is preparing for a 10-mile race.
first 5 miles average speed of 6 miles per hour,
next 3 miles at an average speed of 8 miles per hour,
final 2 miles at an average speed of 10 miles per hour.
..
first leg time = 5/6 hours
second leg time = 3/8 hours
third leg time = 2/10 hours
..
total time = 5/6+3/8+2/10
LCM = 480
5*80 + 3*60+2*48 / 480
400+180+96 / 480
676/480
= 1.41 hours